OpenStudy (anonymous):

A 390 g mass vibrates according to the equation x = 0.38 sin (5.48t) where x is in meters and t is in seconds. Determine the frequency.

jimthompson5910 (jim_thompson5910):

The period of x = A*sin(Bt) is T = 2pi/B In your case, B = 5.48. So the period is T = 2pi/5.48 = 1.14656 The frequency, f, is equal to one over the period. So f = 1/T = 1/1.14656 = 0.87217 So the frequency is 0.87217

OpenStudy (anonymous):

thank you!